OpenStudy (anonymous):

A chemist prepares a solution by dissolving 0.3 mole of CaCl2 in enough water to make 1000 mL of solution. What is the molar concentration of chloride ions? 0.6 M 0.4 M 0.1 M 0.2 M 0.3 M

OpenStudy (anonymous):

@abb0t @rvc

rvc (rvc):

@aaronq

OpenStudy (anonymous):

@abb0t @DoShKa_SyRiA @Frosty_Flow

OpenStudy (abb0t):

Concentration is given by units, \(\sf \large \frac{mol}{L}\) That means you will be dividing moles by Litres of solution.

OpenStudy (anonymous):

But there will be double the number of moles in water of Chloride , right ?

OpenStudy (anonymous):

0.6 / 1 ?

OpenStudy (abb0t):

It is moles of solute per litres of solution.

OpenStudy (anonymous):

But we have to find the molar concentration of Chloride ions . When it will dissolve , there will be two moles of Cl2 for every one mole of CaCl2 . Therefore Moles of Solute = 0.3 * 2 = 0.6 1000 ml = 1L. 0.6 / 1

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

You wil basically have: \(\sf CaCl_{2} \rightarrow Ca^{2+} + 2~Cl^- + H_2O \)

OpenStudy (abb0t):

As you can see, 2(0.3) moles = 0.6 moles. Then divide by litres.

OpenStudy (anonymous):

So 0.6 M , right ?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Had one more question @abb0t , Consider the force of interaction between a cation and an anion. Which of the following does not make the force more attractive? (Recall Coulomb’s law has the formula F = kq1q2/(εr2) *Change the charge on the anion so that it is less negative *Decrease the distance between the ions *Increase the charge on the cation so that it is more positive *Decrease the dielectric constant of the fluid between the ions *All of the other answer choices listed make the force more attractive

OpenStudy (anonymous):

D. The dielectric constant does not affect the force of attraction.